# An Exact Formula for Quantum Entropy Production along Quantum Trajectories

### v1

@@ I Introduction
$d\widehat{\rho}=\mathcal{L}^{\star}(\widehat{\rho})\,dt+\big{(}L\widehat{\rho}+\widehat{\rho}L^{\ast}-\lambda\widehat{\rho}\big{)}\,dI,$ (1)
where $\mathcal{L}(\cdot)=-i[\cdot,H]+\frac{1}{2}[L^{\ast},\cdot]L+\frac{1}{2}L^{\ast}[\cdot,L]$ is Lindblad generator, $H$ is the Hamiltonian of the system, $L$ is the coupling operator of the system to the monitored field, $\mathcal{L}^{\star}$ is the adjoint operator (i.e., $\mathrm{tr}\{\rho\,\mathcal{L}(X)\}\equiv\mathrm{tr}\{\mathcal{L}^{\star}(\rho)\,X\}$ for all trace-class $\rho$ and bounded $X$). The term $I(t)$ is the innovations (statistically a Wiener process) and $\lambda=\mathrm{tr}\{\widehat{\rho}\,(L+L^{\ast})\}$.
We recall that a state is faithful if its density matrix $\rho$ is strictly positive. In such cases we may define its inverse $\frac{1}{\rho}$.
@@
$d\mathsf{S}\big{(}\widehat{\rho}\big{)}=-\mathrm{tr}_{s}\bigg{(}\widehat{\rho}\,\mathcal{L}\ln\widehat{\rho}\bigg{)}\,dt+\Sigma(\widehat{\rho})\,dt-\mathrm{tr}_{s}\bigg{(}\big{(}L\widehat{\rho}+\widehat{\rho}L^{\ast}-\lambda(t)\widehat{\rho}\big{)}\,\ln\widehat{\rho}\bigg{)}\,dI.$ (2)
@@
$\Sigma(\widehat{\rho})$ $=$ $\sum_{k_{1},k_{2}=0}^{\infty}\dfrac{1}{(k_{1}+1)(k_{1}+k_{2}+2)}\binom{k_{1}+k_{2}}{k_{1}}\,$ (3)
×$\times\mathrm{tr}_{s}\bigg{\{}\big{(}-\frac{1}{\widehat{\rho}}\big{)}^{k_{1}+k_{2}+1}\mathrm{ad}^{k_{1}}_{\widehat{\rho}}(L\widehat{\rho}+\widehat{\rho}L^{\ast})\,\mathrm{ad}^{k_{2}}_{\widehat{\rho}}(L\widehat{\rho}+\widehat{\rho}L^{\ast})\bigg{\}}.$
@@
Our proof relies on formulating the problem in discrete time and taking the continuum limit. We begin with an exposition of the discrete time situation.
I.1 State Evolution
I.1.1 Unconditional State Evolution
@@
I.1.2 Conditional State Evolution
@@ I.2 Generic Model
$\approx$ $\rho+\tau(L\rho L^{\ast}+K\rho+\rho K^{\ast}).$
@@
II Entropy Production
Let $\rho$ be a quantum state (density matrix) which we write as $\rho=e^{-\Phi}$ where $\Phi$ is referred to as the surprise. With the standard notation $\langle A\rangle_{\rho}=\mathrm{tr}\big{(}\rho A\big{)}$, the von Neumann entropy of a system in state $\rho$ is given by
@@ III.1 The Unconditional Entropy
III.2 Entropy along a Quantum Trajectory
@@
We will apply the same method to compute the update rule for $\mathsf{S}\big{(}\widehat{\rho_{n}}(\mathbf{y})\big{)}$ for a fixed quantum trajectory generated by $\mathbf{y}$. We shall use the update rule (32) with $\rho=\widehat{\rho}_{n-1}$ and $\varepsilon=\mathcal{L}^{\star}\widehat{\rho}_{n-1}\tau+\bigg{(}L\widehat{\rho}_{n-1}+\widehat{\rho}_{n-1}L^{\ast}-\lambda_{n-1}\widehat{\rho}_{n-1}\bigg{)}\Delta I_{n}$. This time, the bilinear terms in $\varepsilon$ are not negligible so we have to develop Paycha’s formula to second order:
@@
The zeroth term in the Paycha expansion is $\mathsf{S}\big{(}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}$. Here we note that $\mathrm{ad}^{k}_{\rho}(\varepsilon)\approx\mathrm{ad}^{k}_{\rho}(L\rho+\rho L^{\ast})\,\Delta I_{n}$. The first term in the Paycha expansion simplifies because only the $k_{1}=0$ term survives: all the higher terms involve a trace of a commutator so they vanish identically. This leads to
@@
$\mathsf{S}\big{(}\widehat{\rho_{n}}(\mathbf{y})\big{)}$$\approx$ $\mathsf{S}\big{(}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}-\mathrm{tr}_{s}\bigg{\{}\widehat{\rho}_{n-1}(\mathbf{y})\mathcal{L}\ln\widehat{\rho}_{n-1}(\mathbf{y})\big{)}\bigg{\}}\tau-\mathrm{tr}_{s}\bigg{\{}\big{(}L\widehat{\rho}_{n-1}(\mathbf{y})$ (59)
$+$ $\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast}-\lambda_{n-1}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}\ln\widehat{\rho}_{n-1}(\mathbf{y})\bigg{\}}\Delta I_{n}$
@@
×$\times\mathrm{ad}^{k_{1}}_{\widehat{\rho}_{n-1}(\mathbf{y})}(L\widehat{\rho}_{n-1}(\mathbf{y})+\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast})\mathrm{ad}^{k_{2}}_{\widehat{\rho}_{n-1}(\mathbf{y})}(L\widehat{\rho}_{n-1}(\mathbf{y})+\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast})\bigg{\}}\tau.$

### v2

@@ I Introduction
$d\widehat{\rho_{t}}=\mathcal{L}^{\star}(\widehat{\rho}_{t})\,dt+\sqrt{\eta}\,\big{(}L\widehat{\rho}_{t}+\widehat{\rho}_{t}L^{\ast}-\lambda_{t}\widehat{\rho}_{t}\big{)}\,dI(t),$ (1)
where $\mathcal{L}(\cdot)$ is a generator of Lindblad type, $\mathcal{L}(\cdot)=-i[\cdot,H]+\sum_{k}\frac{1}{2}[L_{k}^{\ast},\cdot]L_{k}+\frac{1}{2}L_{k}^{\ast}[\cdot,L_{k}]$, and $\mathcal{L}(\cdot)^{\star}$ is its adjoint (i.e., $\mathrm{tr}\{\rho\,\mathcal{L}(X)\}\equiv\mathrm{tr}\{\mathcal{L}^{\star}(\rho)\,X\}$ for all trace-class $\rho$ and bounded $X$), $H$ is the Hamiltonian of the system, $L$ is the coupling operator of the system to the monitored field (the collapse operator). The term $I(t)$ is the innovations (statistically a Wiener process) and $\lambda_{t}=\mathrm{tr}\{\widehat{\rho}_{t}\,(L+L^{\ast})\}$.
The factor $\eta\in[0,1]$ is the efficiency of the homodyne detector: $\eta=1$ is perfect detection, while $\eta=0$ effectively means that we are left with a dissipative dynamics described by a standard master equation.
The conditional state $\widehat{\rho}_{t}$ is a stochastic process adapted to the filtration generated by the measurements (or equivalently, the innovations process). Our interest shall be in the associated von Neumann entropies $\mathsf{S}\big{(}\widehat{\rho}_{t}\big{)}=-\mathrm{tr}_{S}\big{(}\widehat{\rho}_{t}\ln\widehat{\rho}_{t}\big{)}$ and this will likewise be an adapted stochastic process.
We recall that a state state is pure if and only if it is a rank-one projection; otherwise it is mixed. A mixed state is deemed to be faithful if its density matrix $\rho$ is strictly positive. In such cases we may define its inverse $\frac{1}{\rho}$.
@@
$d\mathsf{S}\big{(}\widehat{\rho}_{t}\big{)}=-\mathrm{tr}_{s}\bigg{(}\widehat{\rho}_{t}\,\mathcal{L}\ln\widehat{\rho}_{t}\bigg{)}\,dt+\Sigma(\widehat{\rho}_{t})\,dt-\mathrm{tr}_{s}\bigg{(}\big{(}L\widehat{\rho}_{t}+\widehat{\rho}_{t}L^{\ast}-\lambda(t)\widehat{\rho}_{t}\big{)}\,\ln\widehat{\rho}_{t}\bigg{)}\,dI.$ (2)
@@
$\Sigma(\widehat{\rho})=\eta\sum_{k_{1},k_{2}=0}^{\infty}\dfrac{1}{(k_{1}+1)(k_{1}+k_{2}+2)}\binom{k_{1}+k_{2}}{k_{1}}\,\,\mathrm{tr}_{s}\bigg{\{}\big{(}-\frac{1}{\widehat{\rho}}\big{)}^{k_{1}+k_{2}+1}\mathrm{ad}^{k_{1}}_{\widehat{\rho}}(L\widehat{\rho}+\widehat{\rho}L^{\ast})\,\mathrm{ad}^{k_{2}}_{\widehat{\rho}}(L\widehat{\rho}+\widehat{\rho}L^{\ast})\bigg{\}}.$ (3)
@@
I.1 Motivation
Let us first remark that the stochastic master equation 1 admits a pure state solution provided the following conditions are satisfied: the system is initially separated from its environment and in a pure state; the environment consists of just the detector (in which case the Lindblad generator is precisely $\mathcal{L}(\cdot)=-i[\cdot,H]+\frac{1}{2}[L^{\ast},\cdot]L+\frac{1}{2}L^{\ast}[\cdot,L]$ where $L$ is the collapse operator); and we have perfect efficiency $(\eta=1)$. Note that this situation generalizes to more than one monitored output field in the obvious way.
In this special situations, we have that $\widehat{\rho}_{t}\equiv|\psi_{t}\rangle\langle\psi_{t}|$ where the pure stochastic state satisfies (see the contribution of J. Kupsch to the standard reference book Decoherence )
$d|\psi_{t}\rangle=-\bigg{(}iH+\frac{1}{2}(L^{\ast}L-\lambda_{t}L+\frac{1}{4}\lambda_{t}^{2})\bigg{)}|\psi_{t}\rangle\,dt+(L-\frac{1}{2}\lambda_{t})\,|\psi_{t}\rangle\,dI(t).$ (4)
The (conditional) von Neumann entropy, of course, vanishes in these cases.
An interesting special situation is where the collapse operator $L$ is self-adjoint. In particular, the question of whether the stochastic states converge to a pure eigenstates of the collapse operator become natural ABBH . For instance, in SvHM an ensemble of $N$ spin-1/2 particles was effectively studied through their total spin vector $\mathbf{J}$ and one performed a continuous homodyne monitoring with collapse operator $M\,J_{z}$ for some coupling constant $M$. Taking $H\equiv 0$ for convenience and assuming perfect detection efficiency, the conditional state will almost surely converge into one of the eigenstates of $J_{z}$, say $|m\rangle$, with probability $p_{m}=\langle m|\rho_{0}|m\rangle$ in accordance with the Born rule. The conditional entropy should then converge to zero while the unconditioned should of course converge to $-\sum_{m}p_{m}\ln p_{m}$. Rather trivially, Theorem 1 would tell us that the rate of change of the conditional entropy would vanish for $\widehat{\rho}$ diagonal in the basis of the collapse operator. (In SvHM , the problem is made more interesting by applying a feedback control through a Hamiltonian $\gamma\,b(t)J_{y}$ where $b(t)$ depends on the measurement record.)
A related problem arises in the theory of quantum feedback networks GouJam09a GouJam09b . One may model a quantum system with proportional feedback of homodyne measurement in a fairly straightforward way using the conditional dynamics Wiseman . However, since one only needs to know that the measurement is being performed and the output fedback without actually knowing the measurement result, one can alternatively describe the average result coherently - see GouJam09b section IV B. Here one is choosing to ignore the measurement record but the fact that a measurement is being made leads to back action - resulting in a unitary model for the system and the detector. This of course changes if one tried to look inside to see what the detector actually detects. An example where this may be of importance is the proposal for autonomous error-correction of a qubit using quantum switches KNPH . Here several switches are coupled to the qubit in a coherent quantum feedback network: again, technically there is no measurement but one could potentially consider inserting detectors so the the qubit and the switches are driven by each others homodyne outputs. The set up in KNPH is specifically designed so that the switches act as dissipators whenever a particular syndrome error occurs to the qubit. While the set up is completely coherent (therefore unitary), one may naturally ask if the energy dissipated by an error correction could leak out as heat thereby telling us that a error-correction had been made. In this example, the switches are acting as regulators giving a quantum analogue of the flyball governor. Related questions about quantum Maxwell’s demon in the setting of superconducting qubits can be found in the work Cottet, et al. QMD
II Discrete Time Models
Our proof of Theorem 1 relies on formulating the problem in discrete time and taking the continuum limit. We begin with an exposition of the discrete time situation.
II.1 State Evolution
II.1.1 Unconditional State Evolution
@@
II.1.2 Conditional State Evolution
@@ II.2 Generic Model
$\mathcal{V}_{s}\rho$ $=$ $\mathrm{tr}_{a}\bigg{(}V(\rho\otimes\sigma)V^{\ast}\bigg{)}\approx\rho+\tau(L\rho L^{\ast}+K\rho+\rho K^{\ast}).$ (26)
@@
III Entropy Production
Let $\rho$ be a quantum state (density matrix) which we write as $\rho=e^{-\Phi}$. With the standard notation $\langle A\rangle_{\rho}=\mathrm{tr}\big{(}\rho A\big{)}$, the von Neumann entropy of a system in state $\rho$ is given by
@@ IV.1 The Unconditional Entropy
IV.2 Entropy along a Quantum Trajectory
@@
We will apply the same method to compute the update rule for $\mathsf{S}\big{(}\widehat{\rho_{n}}(\mathbf{y})\big{)}$ for a fixed quantum trajectory generated by $\mathbf{y}$. We shall use the update rule (33) with $\rho=\widehat{\rho}_{n-1}$ and $\varepsilon=\mathcal{L}^{\star}\widehat{\rho}_{n-1}\tau+\sqrt{\eta}\bigg{(}L\widehat{\rho}_{n-1}+\widehat{\rho}_{n-1}L^{\ast}-\lambda_{n-1}\widehat{\rho}_{n-1}\bigg{)}\Delta I_{n}$. This time, the bilinear terms in $\varepsilon$ are not negligible so we have to develop Paycha’s formula to second order:
@@
The zeroth term in the Paycha expansion is $\mathsf{S}\big{(}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}$. Here we note that $\mathrm{ad}^{k}_{\rho}(\varepsilon)\approx\sqrt{\eta}\mathrm{ad}^{k}_{\rho}(L\rho+\rho L^{\ast})\,\Delta I_{n}$. The first term in the Paycha expansion simplifies because only the $k_{1}=0$ term survives: all the higher terms involve a trace of a commutator so they vanish identically. This leads to
@@
$\mathsf{S}\big{(}\widehat{\rho_{n}}(\mathbf{y})\big{)}$$\approx$ $\mathsf{S}\big{(}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}-\mathrm{tr}_{s}\bigg{\{}\widehat{\rho}_{n-1}(\mathbf{y})\mathcal{L}\ln\widehat{\rho}_{n-1}(\mathbf{y})\big{)}\bigg{\}}\tau$ (60)
$-\mathrm{tr}_{s}\bigg{\{}\big{(}L\widehat{\rho}_{n-1}(\mathbf{y})+\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast}-\lambda_{n-1}\widehat{\rho}_{n-1}(\mathbf{y})\big{)}\ln\widehat{\rho}_{n-1}(\mathbf{y})\bigg{\}}\Delta I_{n}$
@@
×$\times\eta\mathrm{ad}^{k_{1}}_{\widehat{\rho}_{n-1}(\mathbf{y})}(L\widehat{\rho}_{n-1}(\mathbf{y})+\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast})\mathrm{ad}^{k_{2}}_{\widehat{\rho}_{n-1}(\mathbf{y})}(L\widehat{\rho}_{n-1}(\mathbf{y})+\widehat{\rho}_{n-1}(\mathbf{y})L^{\ast})\bigg{\}}\tau.$
@@ V Conclusion
The main result, Theorem 1, gives the expression for the rate of change of entropy of the conditional state along a particular quantum trajectory under the assumption that the state is faithful. We give a series expansion for an entropy production rate $\Sigma$ which involves multiple nest commutators of the collapse operators with the conditional state. Our proof utilized a specific repeated qubit-probe model to approximate the measurement however the continuous limit should be universal.